OpenStudy (anonymous):

How do I integrate this dt=(a+bv)dv initial conditons are V=0 t=0 a and b are constants

OpenStudy (anonymous):

This is the original question. However, all I have done is divided dv on both sides. http://tnypic.net/4a46c.jpg

OpenStudy (anonymous):

t=av+b*(v^2)/2+c put t=0 and v=0 and find out c and put it and find out final ans

OpenStudy (anonymous):

Did you just multiply the bracket by v? Also why did you divide by 2?

OpenStudy (anonymous):

i give the solution after integration

OpenStudy (anonymous):

and integration of vdv=(v^2)/2

OpenStudy (anonymous):

u got it or not reply fast

OpenStudy (anonymous):

I'm an idiot, figured it out. Need more practice on simple indefinite integrals!